Calculate (step by step) the upper bound of a 95% confidencce interval for a sample of (N = 504) with a sample standard deviation of 2.26 and the sample mean of 2.96? Explain the step by step process in detail.
Answer #1

Confidence Interval
\bar{X} \pm t_{\alpha, n-1} S/\sqrt{n}
-1.648 to. ta,n-1 05,504-1
2.96 ± t 0.05,504-1 * 2.26/v504
Upper bound = 2.96+t o.05,504-1 2.26/V504
Upper bound = 3.1259
Upper bound of 95% confidence interval is 3.1259